Esquire Products Inc. expects the following monthly sales:

January $ 47,000 July $ 41,000
February 38,000 August 45,000
March 31,000 September 48,000
April 33,000 October 53,000
May 27,000 November 61,000
June 25,000 December 43,000
Total sales = $492,000

Cash sales are 40 percent in a given month, with the remainder going into accounts receivable. All receivables are collected in the month following the sale. Esquire sells all of its goods for $2 each and produces them for $1 each. Esquire uses level production, and average monthly production is equal to annual production divided by 12.

a. Generate a monthly production and inventory schedule in units. Beginning inventory in January is 31,000 units.
Answer #1

Below is the chart computing units sold per month:

Month Sales SP Units
Jan $     47,000.00 2 23500
Feb $     38,000.00 2 19000
Mar $     31,000.00 2 15500
Apr $     33,000.00 2 16500
May $     27,000.00 2 13500
Jun $     25,000.00 2 12500
Jul $     41,000.00 2 20500
Aug $     45,000.00 2 22500
Sep $     48,000.00 2 24000
Oct $     53,000.00 2 26500
Nov $     61,000.00 2 30500
Dec $     43,000.00 2 21500
Total $ 4,92,000.00 246000

Units sold = Sales value/ Sales price

January Units sold = 47000/ 2 = 23500 units And so on

So total units sold during the year is 246000 which is the sum of monthly unit sold

If opening inventory is 31000 units so those produced during the year were 246000-31000 = 213000

Therefore monthly production will be 213000/12 = 17916.67 or 17917 rounded off

Inventory Schedule is as follows:

Month Opening units Produced Sold Closing units
Jan 31000 17917 23500 25417
Feb 25417 17917 19000 24334
Mar 24334 17917 15500 26751
Apr 26751 17917 16500 28168
May 28168 17917 13500 32585
Jun 32585 17917 12500 38002
Jul 38002 17917 20500 35419
Aug 35419 17917 22500 30836
Sep 30836 17917 24000 24753
Oct 24753 17917 26500 16170
Nov 16170 17917 30500 3587
Dec 3587 17917 21500 4

4 units remaining is only because of the rounding off otherwise there would not be any closing inventory left.

Opening balance + produced - Sold = Closing balance

Opening balance for feb onwards is the closing balance of the previous month